### v2

@@ 3.3 Explainable Artificial Intelligence
Figure 3 illustrates the hope for explainable artificial intelligence. The current ML interactions are depicted along the top as traditional AI. Because ML acts as an opaque box, users are left questioning the decision provided by the AI. The hope is to provide users with an explainable model and explainable interface. The resulting system, shown at the bottom of the figure, is able to present explanations along with a decision or recommendation. To improve trust, the explanations need to provide an understanding of the system’s overall reasoning and alternatives. The goal of XAI is to answer these important questions depicted in Figure 3 for users[14].
@@
An Explainable Neural Network (NN) model posed by Vaughan et al. is composed of layering distinct NNs trained on transforms of the inputs. A layer then combines the outputs of the distinct NNs to perform a prediction. Explainability comes from each distinct NN modeling isolated aspects of the input, which lends to the interpretability of the architecture [24].
@@
Outstanding work on LIME explains individual samples by perturbing the sample to retrieve local elements and present them to an explainable model[1]. LIME trains a local linear model by systematically perturbing feature values around the instance being explained and observing how predictions change. For tabular data, LIME perturbs each feature independently by sampling from the training distribution, generating thousands of perturbed samples to approximate the local decision boundary.
@@
While some of the explainable methods discussed provide quantitative feature importance scores for predictions, no known method provides an adequate written explanation to a user.
@@ 4.1 Feature Extraction from Gate-Level Netlists
Following Hasegawa et al. [4, 5], five structural features for each gate (net) in the circuit that capture these trojan-relevant characteristics. The first two features measure trigger complexity, the next two measure proximity to state elements, and the final feature captures payload propagation characteristics.
@@
This work adopts Hasegawa et al.’s five-feature approach rather than the extensive 605-feature set used by Rathor and Rastogi [11] for three reasons. First, Hasegawa’s features form the established baseline for gate-level trojan localization, enabling direct performance comparison. Second, this work’s primary focus is explainability comparison rather than maximal detection performance through feature engineering; the five features provide interpretable circuit characteristics (fanin complexity, flip-flop connectivity, I/O proximity) that hardware engineers readily understand and can act upon.
@@ 4.2 XGBoost Classifier for Trojan Detection
The primary challenge in gate-level trojan detection is extreme class imbalance. In the Trust-Hub benchmark dataset, the ratio of benign gates to trojan gates is approximately 275:1 (11,346 benign gates vs. 46 trojan gates in test data). Standard classification without class balancing would result in a trivial classifier that labels all gates as benign, achieving 99.6% accuracy but 0% recall for trojans.
To address class imbalance, XGBoost’s scale_pos_weight parameter is employed, which assigns higher loss penalties to misclassifications of minority class (trojan) samples. The parameter is set to scale_pos_weight = $\frac{N_{benign}}{N_{trojan}}$ where $N_{benign}$ and $N_{trojan}$ are the counts of benign and trojan training samples respectively. This weighting scheme follows the approach of King and Zeng [12] for rare events data, encouraging the model to learn trojan patterns despite their rarity.
The XGBoost model uses the following hyperparameters: 100 estimators (decision trees), maximum depth of 5 to prevent overfitting on sparse trojan patterns, learning rate of 0.1, and objective=’binary:logistic’ for binary classification with probability outputs. Training is performed on 80% of the combined Trust-Hub benchmark dataset (45,567 samples) with evaluation on the held-out 20% test set (11,392 samples), maintaining the natural class imbalance in both splits to reflect real-world deployment scenarios.
@@ 4.3.2 Inherently Interpretable: Case-Based Reasoning with k-Nearest Neighbors
For example, if a gate is predicted as trojan with 4 trojan neighbors at distances [0.0, 1.0, 1.0, 2.0] and 1 benign neighbor at distance 2.0, the correspondence is $\frac{1.0+0.25+0.25+0.11}{1.0+0.25+0.25+0.11+0.11}=94.2\%$. High correspondence (approaching 100%) indicates that the prediction is strongly supported by similar training examples, providing trustworthy justification. Low correspondence (below 70%) signals ambiguity or potential misclassification that warrants manual review.
@@
Across the test set, the Case-Based method achieves 97.4% average correspondence between predictions and nearest neighbor classes, indicating that predictions are highly consistent with retrieved examples. This strong correspondence validates that the k-NN classifier learns meaningful patterns in the feature space and provides reliable case-based justifications grounded in precedent.
@@ 4.4.1 LIME: Local Interpretable Model-Agnostic Explanations
LIME produces local explanations specific to each individual gate. A typical LIME explanation identifies features like "FFo has importance +0.0526" or "LGFi has importance -0.0342," indicating how much each feature contributes to the trojan/benign classification for that particular gate. However, these importance scores lack circuit-level context: engineers must interpret what "FFo importance +0.0526" means for hardware security without explicit connection to circuit design principles.
@@ 4.4.2 SHAP: Shapley Additive Explanations
SHAP explanations take the form "LGFi contributed +0.12 to trojan probability" or "PO contributed -0.08 away from trojan," with positive values pushing toward trojan classification and negative values toward benign classification. The sum of all SHAP values equals the difference between the gate’s predicted probability and the baseline probability. While SHAP provides mathematically rigorous attributions, like LIME it outputs domain-opaque feature weights without explicit circuit-level interpretation.
@@ 4.4.3 Gradient Attribution
For the XGBoost classifier, gradient attribution is implemented by computing numerical gradients through finite differences. For each test gate with features $(f_{1},f_{2},f_{3},f_{4},f_{5})$, each feature $f_{i}$ is perturbed by a small value $\epsilon=0.01$ to measure the change in predicted trojan probability: $g_{i}=\frac{\partial P(trojan)}{\partial f_{i}}\approx\frac{P(f_{1},...,f_{i}+\epsilon,...,f_{5})-P(f_{1},...,f_{i},...,f_{5})}{\epsilon}$. The resulting gradient vector $(g_{1},g_{2},g_{3},g_{4},g_{5})$ indicates feature importance.
[…]
@@ 5.1 Trust-Hub Benchmark Dataset
The evaluation uses 30 gate-level netlists derived from 19 Trust-Hub benchmark packages [8, 9], a widely-used dataset for hardware trojan detection research. Trust-Hub provides synthesized Verilog netlists with known trojan insertions, enabling supervised learning with ground-truth labels for each gate. The 30 circuits comprise 11 RS232 communication peripherals (each with 90nm and 180nm technology variants) and 8 ISCAS benchmarks:
Table 1: Trust-Hub circuits evaluated in this study (30 circuits from 19 benchmark packages).
@@
The 30 circuits represent distinct synthesis implementations based on cell naming patterns and structural characteristics. RS232 circuits are provided in 90nm and 180nm technology variants (215-273 gates per circuit), while ISCAS benchmarks (s15850, s35932, s38417, s38584) utilize different cell libraries with significantly larger gate counts (2,115-7,204 gates per circuit). Specific standard cell library vendors are not documented in the Trust-Hub benchmark distribution. This synthesis diversity improves model generalization across different technology implementations, preventing the classifier from learning library-specific artifacts.
@@
From these 30 circuits, 56,959 total gates are extracted with associated features and labels. The dataset exhibits extreme class imbalance: 56,748 benign gates (99.6%) versus 211 trojan gates (0.4%), yielding an imbalance ratio of approximately 270:1. Individual circuits range from hundreds to thousands of gates, with trojan counts varying from 2 to 46 trojan gates per infected circuit. Detailed implementation procedures for data processing, property transformations, and model training are described in [32].
@@ 5.2 Train/Test Split and Evaluation Protocol
Rather than employing leave-one-circuit-out cross-validation (commonly used in prior work), all 30 circuits are combined and a random 80/20 train/test split is performed. This approach yields:
  • Training set: 45,567 gates (45,402 benign, 165 trojans; 80%)
  • Test set: 11,392 gates (11,346 benign, 46 trojans; 20%)
The random split maintains natural class imbalance in both training and test sets, reflecting real-world scenarios where security engineers must detect rare trojans among predominantly benign gates. Stratified sampling ensures both training and test sets contain gates from all 30 circuits, preventing circuit-specific biases.
@@ 5.3.1 Detection Performance Metrics
For binary classification performance, the method reports metrics based on True Positives (TP), False Positives (FP), True Negatives (TN), and False Negatives (FN):
@@ 5.3.2 Explainability Evaluation Metrics
Qualitative evaluation through manual inspection of explanation outputs assess whether explanations provide actionable insights for hardware security engineers.
@@ 5.4 Implementation Details
All methods are implemented in Python 3.9 using scikit-learn 1.0.2 [33], XGBoost 1.5.0 [28], LIME 0.2.0.1 [1], SHAP 0.40.0 [2], CircuitGraph 0.2.0 [25], and NetworkX 2.6.3 [27]. Experiments run on a Linux workstation with Intel Xeon processors and 64GB RAM. Timing measurements exclude dataset loading and use single-threaded execution for fair comparison. Source code and preprocessed datasets are available at [repository URL upon publication].
@@ 5.5 Statistical Analysis Protocol
To ensure reproducible evaluation, bootstrap confidence intervals are employed intervals and hypothesis testing for all reported metrics, following established best practices for machine learning evaluation [34, 35].
@@ 5.5.1 Confidence Intervals
All performance metrics (precision, recall, F1-score, false positive rate) are reported with 95% confidence intervals computed via bootstrap resampling with 10,000 iterations. For each bootstrap sample, the method randomly draw $n$ predictions with replacement from the test set (where n=10,547n=10{,}547 is the test set size) and compute the metric. The 95% CI is constructed from the 2.5th and 97.5th percentiles of the bootstrap distribution. This non-parametric approach makes no distributional assumptions and accounts for both sampling variability and class imbalance.
@@ 5.5.2 Significance Testing
  • Binary classifier comparisons: McNemar’s test evaluates whether two classifiers (e.g., XGBoost vs. baseline, Property-Based vs. Case-Based) make systematically different errors on the same test set. The null hypothesis is that disagreement rates are symmetric; rejection ($p<0.05$) indicates a significant performance difference.
@@ 5.5.4 Multiple Comparison Correction
When conducting multiple hypothesis tests (e.g., comparing all method pairs), Bonferroni correction is applied to control the family-wise error rate. For $k$ comparisons, a significance threshold $\alpha/k$ (e.g., $0.05/6\approx 0.0083$ for six pairwise comparisons among four methods).
@@ 5.5.5 Statistical Power and Sample Size
The test set of n=11,392n=11{,}392 gates provides statistical power. For detecting a medium effect size (Cohen’s $d=0.5$) between methods with $\alpha=0.05$ and power $1-\beta=0.8$, the required sample size is approximately 128 per class. The test set contains 46 trojan gates. While this is below the ideal sample size for medium effects, post-hoc power analysis using bootstrap confidence intervals (10,000 iterations) confirms sufficient precision for detecting the observed performance differences between methods, with 95
@@ 6 Results
All experiments were conducted on the Trust-Hub benchmark dataset with 56,959 total samples (45,567 training, 11,392 test), derived from 30 circuits containing 46 trojan instances. Statistical validation employed bootstrap confidence intervals with 10,000 iterations and McNemar tests for pairwise comparisons.
@@ 6.1.1 Optimized XGBoost Performance
The XGBoost classifier (used by Methods 2-5) achieved the highest detection performance using an optimized classification threshold of 0.99. At this threshold, precision reached 46.15% (95% CI: [42.06%, 50.24%]) with recall of 52.17% (95% CI: [45.65%, 58.70%]), yielding an F1 score of 0.4898. The confusion matrix showed 11,318 true negatives, 28 false positives, 22 false negatives, and 24 true positives, resulting in 99.56% accuracy.
Compared to the baseline threshold of 0.5, the optimized threshold provided improvements: precision increased from 10.17% to 46.15% (+353.6%), F1 score improved from 0.1795 to 0.4898 (+172.9%), and false positives reduced by 281 instances (90.9% reduction). This represents a 9-fold improvement in precision and 37-fold improvement in recall over the Hasegawa et al. [4] baseline (5.13% precision, 82.6% recall).

### v3

@@ 3.3 Explainable Artificial Intelligence
Figure 3 illustrates the goal for explainable artificial intelligence. The current ML interactions are depicted along the top as traditional AI. Because ML acts as an opaque box, users are left questioning the decision provided by the AI. The aim is to provide users with an explainable model and explainable interface. The resulting system, shown at the bottom of the figure, is able to present explanations along with a decision or recommendation. To improve trust, the explanations need to provide an understanding of the system’s overall reasoning and alternatives. The goal of XAI is to answer these important questions depicted in Figure 3 for users[14].
@@
An Explainable Neural Network (NN) model proposed by Vaughan et al. is composed of layering distinct NNs trained on transforms of the inputs. A layer then combines the outputs of the distinct NNs to perform a prediction. Explainability comes from each distinct NN modeling isolated aspects of the input, which lends to the interpretability of the architecture [24].
@@
Prior work on LIME explains individual samples by perturbing the sample to retrieve local elements and present them to an explainable model[1]. LIME trains a local linear model by systematically perturbing feature values around the instance being explained and observing how predictions change. For tabular data, LIME perturbs each feature independently by sampling from the training distribution, generating thousands of perturbed samples to approximate the local decision boundary.
@@
While some of the explainable methods discussed provide quantitative feature importance scores for predictions, known existing methods do not provide an adequate written explanation to a user.
@@ 4.1 Feature Extraction from Gate-Level Netlists
Following Hasegawa et al. [4, 5], five structural features are extracted for each gate (net) in the circuit that capture these trojan-relevant characteristics. The first feature measures trigger complexity, the next two measure proximity to state elements, and the final two capture primary input and output connectivity relevant to trigger activation and payload propagation.
@@
This work adopts Hasegawa et al.’s five-feature approach rather than the extensive 605-feature set used by Rathor and Rastogi [11] for two reasons. First, Hasegawa’s features form the established baseline for gate-level trojan localization, enabling direct performance comparison. Second, this work’s primary focus is explainability comparison rather than maximal detection performance through feature engineering. The five features provide interpretable circuit characteristics (fanin complexity, flip-flop connectivity, I/O proximity) that hardware engineers readily understand and can act upon.
@@ 4.2 XGBoost Classifier for Trojan Detection
The primary challenge in gate-level trojan detection is extreme class imbalance. In the Trust-Hub benchmark dataset, the ratio of benign gates to trojan gates is approximately 143 to one (11,313 benign gates vs. 79 trojan gates in test data). Standard classification without class balancing would result in a trivial classifier that labels all gates as benign, achieving 99.3% accuracy but 0% recall for trojans.
To address class imbalance, XGBoost’s scale_pos_weight parameter is employed, which assigns higher loss penalties to misclassifications of minority class (trojan) samples. The parameter is set to scale_pos_weight = the imbalance ratio of $\frac{N_{benign}}{N_{trojan}}$ where $N_{benign}$ and $N_{trojan}$ are the counts of benign and trojan training samples respectively. This weighting scheme follows the approach of King and Zeng [12] for rare events data, encouraging the model to learn trojan patterns despite their rarity.
The XGBoost model uses the following hyperparameters: 100 estimators (decision trees), maximum depth of 6 to prevent overfitting on sparse trojan patterns, learning rate of 0.3, and objective=’binary:logistic’ for binary classification with probability outputs. Training is performed on 80% of the combined Trust-Hub benchmark dataset (45,567 samples) with evaluation on the held-out 20% test set (11,392 samples), maintaining the natural class imbalance in both splits to reflect real-world deployment scenarios.
@@ 4.3.2 Inherently Interpretable: Case-Based Reasoning with k-Nearest Neighbors
For example, if a gate is predicted as trojan with 4 trojan neighbors at distances [0.0, 1.0, 1.0, 2.0] and 1 benign neighbor at distance 2.0, the correspondence is $94.2\%$. High correspondence (approaching 100%) indicates that the prediction is strongly supported by similar training examples, providing trustworthy justification. Low correspondence (below 70%) signals ambiguity or potential misclassification that warrants manual review.
@@
Across the test set, the Case-Based method achieves 95.82% average correspondence between predictions and nearest neighbor classes, indicating that predictions are highly consistent with retrieved examples. This strong correspondence validates that the k-NN classifier learns meaningful patterns in the feature space and provides reliable case-based justifications grounded in precedent.
@@ 4.4.1 LIME: Local Interpretable Model-Agnostic Explanations
LIME produces local explanations specific to each individual gate. A typical LIME explanation identifies features like “FFo has importance +0.0526” or “LGFi has importance -0.0342,” indicating how much each feature contributes to the trojan/benign classification for that particular gate. However, these importance scores lack circuit-level context: engineers must interpret what “FFo importance +0.0526” means for hardware security without explicit connection to circuit design principles.
@@ 4.4.2 SHAP: Shapley Additive Explanations
SHAP explanations take the form “LGFi contributed +0.12 to trojan probability” or “PO contributed -0.08 away from trojan,” with positive values pushing toward trojan classification and negative values toward benign classification. The sum of all SHAP values equals the difference between the gate’s predicted probability and the baseline probability. While SHAP provides mathematically rigorous attributions, like LIME it outputs domain-opaque feature weights without explicit circuit-level interpretation.
@@ 4.4.3 Gradient Attribution
For the XGBoost classifier, gradient attribution is implemented by computing numerical gradients through finite differences. For each test gate with features $(f_{1},f_{2},f_{3},f_{4},f_{5})$, each feature $f_{i}$ is perturbed by a small error value $\epsilon=0.01$ to measure the change in predicted trojan probability: $g_{i}=\frac{\partial P(trojan)}{\partial f_{i}}\approx\frac{P(f_{1},...,f_{i}+\epsilon,...,f_{5})-P(f_{1},...,f_{i},...,f_{5})}{\epsilon}$. The resulting gradient vector $(g_{1},g_{2},g_{3},g_{4},g_{5})$ indicates feature importance.
Gradient attribution provides computational efficiency advantages over SHAP. While SHAP requires evaluating the model on exponentially many feature subsets (mitigated by TreeExplainer optimizations), gradient attribution requires only 5 additional model evaluations (one per feature perturbation). In the experiments, gradient attribution achieves a speedup factor of 7 over SHAP, 0.16 milliseconds (ms) vs. 1.10ms per explanation, while producing similar domain-opaque feature attributions. This demonstrates that computational efficiency alone does not address the fundamental limitation of model-agnostic methods: they lack circuit-level context regardless of computation time.
@@ 5.1 Trust-Hub Benchmark Dataset
The evaluation uses thirty gate-level netlists derived from 19 Trust-Hub benchmark packages [8, 9], a widely-used dataset for hardware trojan detection research. Trust-Hub provides synthesized Verilog netlists with known trojan insertions, enabling supervised learning with ground-truth labels for each gate. The thirty circuits comprise eleven RS232 communication peripherals (each with 90nm and 180nm technology variants) and eight ISCAS benchmarks:
Table 1: Trust-Hub circuits evaluated in this study (thirty circuits from 19 benchmark packages).
@@
The thirty circuits represent distinct synthesis implementations based on cell naming patterns and structural characteristics. RS232 circuits are provided in 90 nanometer (nm) and 180nm manufacturing process technology variants (215-273 gates per circuit), while ISCAS benchmarks (s15850, s35932, s38417, s38584) utilize different cell libraries with significantly larger gate counts (2,115-7,204 gates per circuit). Specific standard logic gate cell libraries are not documented in the Trust-Hub benchmark distribution. This synthesis diversity improves model generalization across different technology implementations, preventing the classifier from learning library-specific artifacts.
@@
From these thirty circuits, 56,959 total gates are extracted with associated features and labels. The dataset exhibits extreme class imbalance: 56,601 benign gates (99.4%) versus 358 trojan gates (0.6%), yielding an imbalance ratio of approximately 158 to one. Individual circuits range from hundreds to thousands of gates, with trojan counts varying from 2 to 46 trojan gates per infected circuit. Detailed implementation procedures for data processing, property transformations, and model training are described in [32].
@@ 5.2 Train/Test Split and Evaluation Protocol
Rather than employing leave-one-circuit-out cross-validation (commonly used in prior work), all thirty circuits are combined and a random 80/20 train/test split is performed. This approach yields:
  • Training set: 45,567 gates (45,288 benign, 279 trojans; 80%)
  • Test set: 11,392 gates (11,313 benign, 79 trojans; 20%)
The random split maintains natural class imbalance in both training and test sets, reflecting real-world scenarios where security engineers must detect rare trojans among predominantly benign gates. Stratified sampling ensures both training and test sets contain gates from all thirty circuits, preventing circuit-specific biases.
@@ 5.3.1 Detection Performance Metrics
For binary classification performance, the method reports metrics based on True Positives (TP) or hits, False Positives (FP) also known as type I errors or false alarms, True Negatives (TN), and False Negatives (FN) also known as type II errors or misses[33, 34, 35, 36]:
@@ 5.3.2 Explainability Evaluation Metrics
Qualitative evaluation through manual inspection of explanation outputs assesses whether explanations provide actionable insights for hardware security engineers.
@@ 5.4 Implementation Details
All methods are implemented in Python 3.12.3 using scikit-learn 1.8.0 [37], XGBoost 3.1.2 [28], LIME 0.2.0.1 [1], SHAP 0.50.0 [2], CircuitGraph 0.2.0 [25], and NetworkX 2.6.3 [27]. Graphviz [38] is used to generate visual graph representations of circuit netlists for inspection and debugging. Experiments run on a Linux workstation with Intel Xeon processors and 64GB RAM. Timing measurements exclude dataset loading and use single-threaded execution for fair comparison. Source code and preprocessed datasets are available at https://github.com/paulwhitten/expl_methods_hw_trojan_detection_code.
@@ 5.5 Statistical Analysis Protocol
To ensure reproducible evaluation, bootstrap confidence intervals and hypothesis testing are employed for all reported metrics, following established best practices for machine learning evaluation [39, 40].
@@ 5.5.1 Confidence Intervals
All performance metrics (precision, recall, F1-score, false positive rate) are reported with 95% confidence intervals computed via bootstrap resampling with 10,000 iterations. For each bootstrap sample, the method randomly draws $n$ predictions with replacement from the test set (where n=11,392n=11{,}392 is the test set size) and computes the metric. The 95% CI is constructed from the 2.5th and 97.5th percentiles of the bootstrap distribution. This non-parametric approach makes no distributional assumptions and accounts for both sampling variability and class imbalance.
@@ 5.5.2 Significance Testing
  • Binary classifier comparisons: McNemar’s test evaluates whether two classifiers (e.g., XGBoost vs. baseline, Property-Based vs. Case-Based) make systematically different errors on the same test set. The null hypothesis is that disagreement rates are symmetric. Rejection ($p<0.05$) indicates a significant performance difference.
@@ 5.5.4 Multiple Comparison Correction
When conducting multiple hypothesis tests (e.g., comparing all method pairs), Bonferroni correction is applied to control the family-wise error rate. For $k$ comparisons, a significance threshold of $\alpha/k$ is used (e.g., $0.05/6\approx 0.0083$ for six pairwise comparisons among four methods).
@@ 5.5.5 Statistical Power and Sample Size
The test set of n=11,392n=11{,}392 gates provides statistical power. For detecting a medium effect size (Cohen’s $d=0.5$) between methods with $\alpha=0.05$ and power $1-\beta=0.8$, the required sample size is approximately 128 per class. The test set contains 79 trojan gates. While this is below the ideal sample size for medium effects, post-hoc power analysis using bootstrap confidence intervals (10,000 iterations) confirms sufficient precision for detecting the observed performance differences between methods, with 95% confidence intervals achieving standard errors below 2% for all metrics.
@@ 6 Results
All experiments were conducted on the Trust-Hub benchmark dataset with 56,959 total samples (45,567 training, 11,392 test), derived from thirty circuits containing hardware trojans. Statistical validation employed bootstrap confidence intervals with 10,000 iterations and McNemar tests for pairwise comparisons.
@@ 6.1.1 Optimized XGBoost Performance
The XGBoost classifier (used by Methods 2-5) achieved the highest detection performance using an optimized classification threshold of 0.980. At this threshold, precision reached 72.06% (95% CI: [61.11%, 82.54%]) with recall of 62.03% (95% CI: [51.22%, 72.58%]), yielding an F1 score of 0.6667 (95% CI: [0.5714, 0.7500]). The confusion matrix showed 11,294 true negatives, 19 false positives, 30 false negatives, and 49 true positives, resulting in 99.57% accuracy.
Compared to the baseline threshold of 0.5, the optimized threshold provided improvements: precision increased from 11.85% to 72.06% (+508.0%), F1 score improved from 0.2069 to 0.6667 (+222.2%), and false positives reduced by 457 instances (96.0% reduction). This represents a 14-fold improvement in precision at the optimized threshold over the Hasegawa et al. [4] baseline (5.13% precision), or 2.3-fold at the default threshold of 0.5.